OpenStudy (anonymous):

Simplify. 7-1(-1 is exponent)+7-2(same)/ 7-3

OpenStudy (aum):

\[\large \frac{7^{-1}+7^{-2}}{7^{-3}} = (7^{-1}+7^{-2}) * 7^3 = \\ \large 7^{-1}*7^3 + 7^{-2}*7^3 = 7^2+7^1 = 49+7 = 56 \]

OpenStudy (anonymous):

What happnes to the 7 3?

OpenStudy (aum):

\(\large 7^{-3}\) in the denominator will become \(\large 7^{3}\) when brought to the numerator.

OpenStudy (aum):

\[ \large \frac{1}{a^{-m}} = a^{m} \\ \large \frac{1}{a^m} = a^{-m} \\ \]

OpenStudy (anonymous):

no but what happnes to it it just sisappers soesnt do anything ot the other 7s

OpenStudy (anonymous):

disappears

OpenStudy (aum):

\[ \large \frac{7^{-1}+7^{-2}}{7^{-3}} = (7^{-1}+7^{-2}) * 7^3 = 7^{-1}*7^3 + 7^{-2}*7^3 \]Did you follow up to this point?

OpenStudy (anonymous):

no u distribute but it doesn show anythin in amswer

OpenStudy (aum):

I am asking you if you followed up to that point. If you did not follow up to that point tell me where you are having doubts.

OpenStudy (anonymous):

O just realised that u dont do anything with the sevens just the expoents i think i got it now thhnx

OpenStudy (aum):

\[ \large 7^{-1} * 7^3 = 7^{-1+3} = 7^{2} = 49\\ 7^{-2} * 7^3 = 7^{-2+3} = 7^{1} = 7 \\ 49 + 7 = 56 \]